OpenStudy (anonymous):

Write an equation for the line that is parallel to the given line and passes through the given point. y = x + 8; (4, 10)

hartnn (hartnn):

can u identify the slope of given line y=x+8 ?

hartnn (hartnn):

the slope is 1, if u compare it with y=mx+c , m=1 so your line equation now is y=x+c now just put x=4, y=10 to get c.

OpenStudy (anonymous):

and that would b ??

hartnn (hartnn):

10= 4+c so c=6 so your line equation is y=x+6. ok ?

OpenStudy (anonymous):

y = x – 6

hartnn (hartnn):

nopes. y=x+6

OpenStudy (anonymous):

hmmmmmmmmmmm o i c :P thanks !!!

hartnn (hartnn):

welcome :)